a definite amount of gaseous hydrocarbon (having carbon atoms less than 5) was burnt with sufficient amount of oxygen.the volume of all reactants was 600 ml,after the explosion the volume of the products [CO2(g) and water(g)]was found to be 700 ml under the similar conditions .what is the molecular formula of the compound?

Grade:

1 Answers

Rahul askIITiansExpert.IITR
13 years ago
it is a standard standard hydrocarbon combustion
 
CxHy +(x+y/4)CO2= xCO2 +(y/2) H2O

for detm. of formula of hydrocarbon from which the questions usually come--
4(p/m-1)=y
mx=n     from x,y detm. of formula is done.
where n=vol. of CO2      p=diff. in volume      m=vol. of hydrocarbon
I think you wil get the answer.